OpenStudy (anonymous):

When you subtract 4x to the 4 from 2x to the fourth why don't the x to the 4 change?

TheSmartOne (thesmartone):

Can you give us the full exact question, so we know what you are referring to?

TheSmartOne (thesmartone):

@SelenaPhillips

OpenStudy (anonymous):

4x to the 4 power minus 2x to the 4 power is what I am having trouble with.

TheSmartOne (thesmartone):

Ah, so \(\sf\Large (4x)^4 - (2x)^4 = ?\) That is your question, right?

OpenStudy (anonymous):

Yes

TheSmartOne (thesmartone):

I guess your using the app, right? so you won't be able to see that, so: (4x)^4 - (2x)^4 = ?

TheSmartOne (thesmartone):

I guess your using the app, right? so you won't be able to see that, so: (4x)^4 - (2x)^4 = ?